OpenStudy (anonymous):

Solve. x-square root4-3x=-8

OpenStudy (anonymous):

x+8=sqrt(4-3x) x^2+16x+64=4-3x x^2+19x+60=0 (x+15)(x+4)=0 x--15,-4

OpenStudy (anonymous):

Confused about the second line. Can you explain.

OpenStudy (anonymous):

i did square of both side

OpenStudy (anonymous):

How did you get 64 and after squaring, why does the 4-3x remain?